Nominal GDP in Zambia has grown rapidly in recent years. But historically, Zambia is a country that has struggled with inflation rates. The table below gives statistics for Zambia in a recent year.

Nominal GDP in Zambia has grown rapidly in recent years. But historically, Zambia is a country that has struggled with inflation rates. The table below gives statistics for Zambia in a recent year.
 

Nominal GDP growth rateGDP deflator growth ratePopulation growth rate
 14.8%7.4%1.1%



What was the rate of economic growth for Zambia?     % (Round your answer to the nearest tenth.)
 

Part 2   (1 point)

If Zambia continues to grow at its current rate of economic growth, how long will it take to double the level of per capita real GDP? Round your answer to the nearest tenth.
     years 
Answer #1

Zambia’s economic growth rate was 14.8% - 7.4% - 1.1% = 6.3%. 


Part 2. At the current rate of growth, it will take Zambia 11.1 years to double its per capita real GDP. Using the rule of 70 -> 70/6.3 =11.1 

answered by: Maggie1920
Answer #2

A]


Here, Nominal GDP growth rate = 14.8%


GDP deflator growth rate = 7.4%


Population growth rate = 1.1%


The rate of economic growth is calculated by the growth rate in real GDP. Therefore, the economic growth rate for nation Z can be calculated as follows:


Real GDP growth rate = Nominal GDP growth rate - GDP Deflator growth rate


- Population growth rate


= (14.8 – 7.4 – 1.1)% = 6.3%


Therefore, the economic growth rate of nation Z is 6.3%


B]


“Rule of 70' is the rule to find out how many years it takes for an economy to double its GDP. It can be calculated as follows:


Number of years taken to double the GDP = 70 / Economic growth rate


Since the economic growth rate is 6.3%, therefore,


Number of years taken to double the GDP = 70 / 6.3


= 11.11years = 10years (nearest tenth)


Therefore, it will take 10 years for the nation Z to double its per capita real GDP.


answered by: Olivier